(3) How many years would it take for you to double your initial deposit? The interest rate is 4%/ year, compounded annually. You make a deposit only once in the beginning. You may round the final answer to the nearest hundredth. For example, just write down 3.41 years if the true answer were 3.4123. (Hint: Let SC denote the initial deposit. Then, the future value will be $2C someday. Find how many years you should wait until such day comes finally, using the future value formula. You can do it by trial and error or by using log(2) and log(1.04). If you dont know the logarithm, just do it by trial and error. When you do it by trial and error, you can replace SC by any dollar amounts (say S1), and your answer will be still the same because SC will be cancelled out during calculation anyway.) 1. 17.67years 2. 18.67years 3. 19.67years 4. 20.67years 5. 21.67years
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text General aCopy в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Format Paste Sort &Find & Format Painter Formatting as Table Styles2 Clear Clipboard AY172 AR Font Alignment Number Styles Cells Edting AS AT AU AV AW AX AY AZ BA BC BD 167 168 169 170 PV(1+0.04)An 1(1+0.04)An (1+0.04)An FV= 172 173 174 175 176 BY TRIAL & ERROR, WE KNOW THAT n IS BETWEEN 17 &18 AT n-17 YEARS, FV- (1.04)A17- 1.9479 AT n = 18 YEARS, FV = ( 1.04)^ 18-2.0258 LOWER VALUE HIGHER VALUE REQUIRED VALUE = 2 178 179 180 181 182 183 184 185 FORMULA OF CALCULATING n n NO OF YEARS AT HIGHER VALUE- (HIGHER VALUE-REQUIRED VALUE)/(HIGHER VALUE-LOWER VALUE) n 18-(2.0258-2)/(2.0258-1.9479) 17.67 YEARS | CALCULATOR AN OPTIONS fv. ann CAP STRU VALUE BOX ! DILUTION DOLLAR COST AVG MORTGAGE.. EXPO ( Sheet1 . Sheet2. Shell- 福 130% 03:44 21-01-2019